OpenStudy (anonymous):

radical 5 plus radical 2

OpenStudy (anonymous):

pretty sure that's as far as you can go

OpenStudy (anonymous):

\[\sqrt{5} + \sqrt{2}\]

OpenStudy (anonymous):

whats the answe?

OpenStudy (unklerhaukus):

are you looking for an approximation?

OpenStudy (anonymous):

\[\approx 3.65\]

OpenStudy (anonymous):

in radical form

OpenStudy (anonymous):

not decimal

OpenStudy (unklerhaukus):

i dont think you can simplify \(\sqrt 5+\sqrt 2\)

OpenStudy (anonymous):

@UnkleRhaukus , correct :)

OpenStudy (anonymous):

You just leave the answer as an addition of radicals.

OpenStudy (anonymous):

ok so this is the question

OpenStudy (anonymous):

(sq root of 5 + sq root of 2) ( sq root of 5 - sq root of 2)

OpenStudy (anonymous):

\[\sqrt(5+\sqrt(2)) \sqrt(5-\sqrt(2))\] \[\sqrt((5-\sqrt(2)) (5+\sqrt(2)))\]

OpenStudy (anonymous):

so... it would be \[\sqrt(23)\]

OpenStudy (anonymous):

thank you soooo mucchhh but how did u get it? :D

OpenStudy (anonymous):

No that's incorrect.

OpenStudy (anonymous):

\[(\sqrt{5}+\sqrt{3})(\sqrt{5}-\sqrt{3})=(\sqrt{5}^{2}-\sqrt{5}\sqrt{3}+\sqrt{5}\sqrt{3}-\sqrt{3}^{2})\]

OpenStudy (anonymous):

sorry replace all the "3" with "2"

OpenStudy (anonymous):

\[=(5-2)=3\]

OpenStudy (anonymous):

that' s your answer

OpenStudy (anonymous):

I will be back

OpenStudy (anonymous):

Actually i got 2 for what you wrote.

OpenStudy (anonymous):

^for ur second post